Refer to the following partial one-factor ANOVA results from excel (some information is missing. And you...

  1. Refer to the following partial one-factor ANOVA results from excel (some information is missing. And you need to work it out in the questions below.) Now, the F statistic is equal to:

Source of Variation

Sum of Squares

Degrees of Freedom

Mean Square

F statistic

Between Groups

210.2788

Within Groups

1483

74.15

Total

2113.833

  1. 4.79             
  2. 3.56
  3. 1.15
  4. 2.84
  1. Referring to the table in question 1. The sum of squares for between groups variation is:
  1. 129.99
  2. 630.83
  3. 1233.4
  4. We cannot tell from given information
  1. Referring to the table in question 1, the Degree of Freedom for between groups variation is (hint: use the results from Q2 and the mean square of between groups):
    1. 3
    2. 4
    3. 5
    4. 6
Answer #1

Using above ANOVA result we get,

i) the F statistic is,

F = MSBetween / MSWithin = 210.2788 / 74.15 = 2.84

Answer: d) 2.84

ii) The sum of squares for between groups variation ( SSBetween ) is,

SSBetween = SSTotal - SSWithin = 2113.833 - 1483 = 630.833

Answer: b) 630.83

iii) Degrees of Freedom for between groups variation ( DFBetween ) is,

DFBetween = SSBetween / MSBetween = 630.83 / 210.2788 = 2.999

DFBetween = 3

Answer: a) 3

DFWithin = SSWithin / MSWithin = 1483 / 74.15 = 20

DFTotal = DFBetween + DFWithin = 3 + 20 = 23